OpenStudy (anonymous):

Select the equation of a line that is perpendicular to the line on the graph and passes through the point (3, 2).

OpenStudy (alexandervonhumboldt2):

perpendicular lines have negative reciprocal slopes

OpenStudy (anonymous):

y = 1 over 3 x + 3 y = -1 over 3x + 3 y = 3 x + 2 y = - 1 over 3 x + 2 are the choices h

OpenStudy (anonymous):

Maybe B?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

cam you do 1 more please

OpenStudy (alexandervonhumboldt2):

yeah